# A Landmark-Aware Visual Navigation Dataset

### v1

@@ Abstract.
Map representation learned by expert demonstrations has shown promising research value. However, recent advancements in the visual navigation field face challenges due to the lack of human datasets in the real world for efficient supervised representation learning of the environments. We present a Landmark-Aware Visual Navigation (LAVN) dataset to allow for supervised learning of human-centric exploration policies and map building. We collect RGB observation and human point-click pairs as a human annotator explores virtual and real-world environments with the goal of full coverage exploration of the space. The human annotators also provide distinct landmark examples along each trajectory, which we intuit will simplify the task of map or graph building and localization. These human point-clicks serve as direct supervision for waypoint prediction when learning to explore in environments. Our dataset covers a wide spectrum of scenes, including rooms in indoor environments, as well as walkways outdoors. Dataset is available at DOI: 10.5281/zenodo.10608067.
@@
Human-in-the-Loop, Visual Navigation, Map Representation, Graph Representatin, Dataset
ccs: Information systems Multimedia databasesccs: Computing methodologies Machine learningccs: Human-centered computing Interactive systems and tools
1. Introduction
@@
Robotic navigation remains a research problem of interest as advances in AI make mobile intelligent agents, such as delivery bots, more feasible. Many approaches have been proposed from SLAM (Mur-Artal et al., 2015; Mur-Artal and Tardos, 2017; Campos et al., 2021), to graph-based methods (Bavle et al., 2023, 2022; Kim et al., 2023; He et al., 2023), to reinforcement learning (RL) (Zhu et al., 2017) to enable robots to effectively move around a space in a goal-oriented manner. However, these methods suffer from high computational complexity, high memory usage, and low sample efficiency respectively. Visual navigation (Ramakrishnan et al., 2022; Morin et al., 2023; Hahn et al., 2021) methods circumvent several of these challenges by using neural networks to implicitly learn scene geometry and an optimal navigation policy without necessitating direct environment interaction.
@@
Figure 1. We propose a novel Landmark-Aware Visual Navigation (LAVN) dataset for map representation research. LAVN provides unique landmark information (in red) in specific scenes to aid model navigation and exploration performance and map representation learning. In this figure, a human expert in Room A (in blue) provides a video-based training trajectory (in green lines) for navigation given a goal image (green box). Rooms are depicted in different colors and represented with nodes. Our dataset facilitates topological map representation learning by injecting novel landmarks into a robot’s observation.
@@
Typically, visual navigation methods are evaluated with the image goal task, where an agent is given a goal image of some target and tasked with finding and reaching it in the environment. The success of state of the art (SOTA) methods at this task comes down to an agent’s ability to efficiently explore its environment. Several works focus on explicitly learning optimal exploration policies (Chaplot et al., 2020b; Ye et al., 2021; Chaplot et al., 2021, 2020a), but this is not the case for the majority of SOTA methods. They either abstract exploration out of the problem all together by assuming that the environment has been seen by the agent prior to test time (Morin et al., 2023), assume that their networks will implicitly learn an adequate exploration strategy from the data due to the specific task structure (Krantz et al., 2021), or explicitly define the exploration strategy for the agent by hard coding it into the architecture (Hahn et al., 2021).
Dataset Human Trajectory Sim Real VQA Large Scale Semantics Landmarks
Demonstrations
Habitat-Web (Ramrakhya et al., 2022)
RUGD (Wigness et al., [n. d.])
RealEstate20k (Zhou et al., 2018)
EmbodiedQA (Das et al., 2018)
@@
Table 1. Summary of dataset characteristics for visual navigation. Different from existing datasets, LAVN provides video-based trajectories guided by humans in both simulated and real-world scenarios. We are the only dataset to provide labels for helpful navigational landmarks. These landmarks introduce little overhead to an agent’s learning scheme, as they do not necessitate the interpretation of the additional text modality as in VQA, or heavy semantic understanding of the environment. We are also the only dataset to provide human trajectory demonstrations in both real world and simulated environments. While we do not provide millions of data points, we allow for supervised model training, which removes the need for copious amounts of data.
Of those methods that learn to explore, several niches have arisen such as adding an exploration term to RL reward functions with (Ye et al., 2021) and without (Chen et al., 2019) behavioral cloning, building metric maps or graphs with (Chaplot et al., 2021) and without (Chaplot et al., 2020c) semantics, and learning to propose exploration waypoints that provide helpful subgoals during navigation for local policies (Chaplot et al., 2020a). These methods suffer from similar downsides as those mentioned previously with the added potential for also requiring full ground truth semantic labels of an environment, which can be extremely costly to acquire. Instead, a much simpler approach would be to explicitly learn good exploration techniques in a human-centric manner. Learning human-centric exploration strategies in a supervised manner would allow for much simpler networks to effectively solve navigation problems, thus reducing the computational complexity, high memory usage, and low sample efficiency issues of SOTA. However, there is very limited data that would allow for this kind of learning.
In this work, we present a Landmark-Aware Visual Navigation (LAVN) dataset [^1] Dataset is available at https://zenodo.org/records/10608067. of human-guided environment exploration videos collected from the photo-realistic Habitat virtual environment as well as several real world environments. We collect RGBD image observations and human point click pairs as a human guides an agent in exploring each environment. These point clicks represent subgoals that guide the agent incrementally though the space and can be used for supervised learning of human-centric waypoints for environment exploration. Additionally, we expect that having landmarks when exploring has the potential to improve current SOTA performance on map-building and localization tasks inherent in navigation. To this end, the human annotator also provides examples for helpful navigation landmarks, which also allows for supervised training of landmark prediction networks. In this way, it becomes possible to directly train a network to explore an environment in a human-centric manner under multiple SOTA paradigms.
This paper is organized as follows: we summarize our dataset’s key differences from existing works in Section 2. Then, we present the details of data acquisition in both the virtual (Section 3) and real-world (Section 4) environments. The details of the dataset format will be described in Section 5. We conclude our work in Section 6.
2. Related Work
2.1. Map Representation for Navigation
The success of navigation relies on an effective learning and understanding of the environment. Two main approaches for this are map representation learning in geometry-based (Mur-Artal et al., 2015; Mur-Artal and Tardos, 2017; Campos et al., 2021) and learning-based (Chaplot et al., 2020b; Zhu et al., 2017; Morin et al., 2023; Shah et al., 2021; Vezhnevets et al., 2017) categories. However, the former requires careful and robust hand-engineered feature design and requires intensive computation (e.g. keyframes in SLAM). The advancements of Deep RL in the second category are partially attributed to the emergence of simulators designed for photorealistic simulation, including Habitat AI (Szot et al., 2021; Savva et al., 2019) with datasets Gibson (Xia et al., 2018) and Matterport (Chang et al., 2017). While these simulators enable large scale model training, they are also expensive to build, can run slowly enough to mitigate their ability to produce large numbers of training samples, and introduce the sim2real problem when deploying models in the real world. To combat this, methods like NRNS (Hahn et al., 2021) demonstrate the effectiveness of using expert videos to train robots without interacting with environments.
2.2. Human-Robot Interaction Datasets
There is a large body of datasets or benchmarks for Human-Robot Interaction, ranging from industrial (Iodice et al., 2022) to social (Karnan et al., 2022; Rudenko et al., 2020; Day and Karamouzas, 2023; Wan et al., 2022; Anvari and Wurdemann, 2020; Schreiter et al., 2022; Shu et al., 2016; Chen et al., 2022) environments. As one of the fundamental tasks for robots, researchers have collected datasets for learning visual navigation with socially compliant behaviors (Karnan et al., 2022; Rudenko et al., 2020; Day and Karamouzas, 2023; Anvari and Wurdemann, 2020; Schreiter et al., 2022; Shu et al., 2016; Chen et al., 2022). Apart from learning from other robots collectively (Shah et al., 2023), robotic visual navigation can benefit from human feedback via visual cues (Wigness et al., [n. d.]), language (Thomason et al., 2020), gesture (Cao, 2018), multimodal instructions (Schreiter et al., 2022; Shrestha et al., 2023), or demonstrations (Ramrakhya et al., 2022).
Our work extends the line of efficient map representation using topological graphs (He et al., 2023; Kim et al., 2023). Similar to how a human (Chrastil and Warren, 2014) can quickly learn to navigate in a new environment by implicitly treating landmarks as nodes in their mental graph (Chrastil and Warren, 2014), we hypothesize that a robot can acquire the ability to navigate in novel places efficiently by intelligently injecting landmarks into its own topological graph. However, little attention has been paid to such datasets, resulting in a lack of data that includes landmark information for this research in the literature. Our dataset hereby aims to fill this gap by providing a series of human-expert-annotated video with landmarks both in virtual and real-world datasets, including indoor and outdoor scenes. We summarize the key characteristics of recent datasets in Table 1.
@@ 2.2. Human-Robot Interaction Datasets
Figure 2. Samples of RGB images in our dataset captured from an ego-centric camera. Video frames start from left to right. Landmarks are visualized with red dots. Samples in the virtual dataset (1st and 2nd rows in blue) are visually realistic compared to the real-world observations (3rd and 4th rows in green), which makes landmark data acquisition scalable to serve real-world research purposes.
3. Virtual Dataset
@@
The main objective of the data collection is to provide human navigation demonstrations for robots in a real-world environment. As video has been proven to be an effective way to teach robots to accomplish different tasks (Hahn et al., 2021), we hereby collect trajectories in video format. Below, we detail the process of data collection for the virtual dataset.
3.1. Environments
@@
We provide human-click annotations for a selection of the Matterport (Chang et al., 2017) and Gibson environments (Xia et al., 2018) in Habitat-Sim (Szot et al., 2021; Savva et al., 2019) to facilitate virtual environment exploration. We provide a full list of annotated environments with the data. Habitat is already used to train many visual navigation methods, so having in-domain, landmark-aware data will allow for direct extensions of these methods for greater ease of model comparison. Additionally, it allows for greater data collection volume and lower data processing time.
3.2. Data Collection and Representation
@@
To start, the human navigator is placed at a random, navigable point in an environment, and tasked with visiting as many rooms as possible while identifying distinct landmarks in each room. The current first person observation of the environment is shown on a screen, and the human moves through the environment by left clicking on this image towards the direction they would like to move. There are three discrete action choices for moving around in the environment: moving forward 0.25m, turning left $15^{\circ}$, and turning right $15^{\circ}$. To specify a landmark, the navigator right clicks on the observation image over the object or region of space they choose. We provide a human point-click supervisory signal for each time step of the trajectory, but these signals could easily be masked to generate data to train temporally abstracted waypoint predictors.
@@
A graph is built as the navigator moves through the space, where nodes are added each time an action is taken. Each node contains one RGB observation, one depth image, the human navigational point click, and the ground truth odometry (including ground truth location and orientation). Edges are drawn between consecutive nodes only, however it is possible to do loop closure on the graph using the odometry provided. Each graph also contains a sequential list of the actions taken to move through the environment. A list of pairs of landmark point-click image coordinates paired and their corresponding nodes is also provided in the graph so that landmarks may be easily identified in the scene. Each human-guided trajectory is terminated after 500 actions or when the annotator has navigated through the entirety of the space, whichever comes first.
@@
Figure 3. Samples of RGB images in the real world environments with landmarks annotated by red dots. Trajectories are recorded in comprehensive environments both indoors and outdoors.
4. Real World Dataset
In this section, we detail the process of data collection for the real-world dataset. We aim to make the data collection and representation steps between the virtual and real-world datasets as similar as possible.
4.1. Environments
@@
4.2. Data Collection and Representation
@@
In order to collect consistent data between the real-world data and simulation, our human annotator is asked to follow the same protocol as they did in the virtual environment described in Section 3.2. During navigation, the annotator holds a camera (iPhone13Mini (iph, 2021)’s built-in camera) naturally around 1.5m high off the ground facing forwards. Unlike in simulation where agents can move with exact distances or turning angles, real-world environments require that each action taken by the human navigator is inherently inexact. Therefore, the navigator is instructed to follow the same discrete action choices by its own judgement. Specifically, the expert takes a picture (saved in JPEG) in every step with around 0.7 meters, or in every left or right turn in roughly $15^{\circ}$. This also presents real-world data with complex noise or perturbations. Trajectories are recorded in ten diverse scenes in both indoor and outdoor environments shown in Fig. 3.
@@
Resolution in the original images was 4032x3024. To promote consistency between the virtual and real-world data, the images are resized to be the same resolution of 640x480 as in the simulation and saved in JPEG format with quality $85\%$. Landmarks are annotated offline after the videos are saved. Samples from our dataset are visualized in Fig. 2. Both virtual and real-world images are photorealistic and can benefit downstream research in the real world. Our proposed dataset provides a unique benchmark for both learning and evaluation to serve the visual navigation community. Our dataset statistics are summarized in Table 2.
Dataset ##\##Traj. Tot. ##\##Frames Tot. ##\##Virt. ##\##Real World
##\##Landmarks Envs Envs
LAVN 310 103,998 14,281 300 10
Table 2. Statistics of our dataset. We detail the number of trajectories, total number of frames per trajectory, landmarks, virtual environments, and real world environments included in the data.
5. File Formatting
For both datasets, the graph created during navigation is saved as a json file containing edge, action, odometry, point-click, and image path information for each node. The RGB observation and depth images for the virtual dataset are saved as png’s of size 480x640. For the real world dataset, the RGB observations and depth images are saved as jpg’s of the same size.
The overall file structure of the dataset is separated into different folders for the virtual and realworld datasets. Within the virtual dataset folder, there are different folders for the Gibson and Matterport data, where each trajectory has it’s own folder within one of these and is labeled with it’s environment name. We provide tools for video playback, reading in the graphs, and processing the images with the data.
6. Conclusion

### v2

@@ Abstract
Map representations learned by expert demonstrations have shown promising research value. However, the field of visual navigation still faces challenges due to the lack of real-world human-navigation datasets that can support efficient, supervised, representation learning of environments. We present a Landmark-Aware Visual Navigation (LAVN) dataset to allow for supervised learning of human-centric exploration policies and map building. We collect RGBD observation and human point-click pairs as a human annotator explores virtual and real-world environments with the goal of full coverage exploration of the space. The human annotators also provide distinct landmark examples along each trajectory, which we intuit will simplify the task of map or graph building and localization. These human point-clicks serve as direct supervision for waypoint prediction when learning to explore in environments. Our dataset covers a wide spectrum of scenes, including rooms in indoor environments, as well as walkways outdoors. We release our dataset with detailed documentation at https://huggingface.co/datasets/visnavdataset/lavn (DOI: 10.57967/hf/2386) and a plan for long-term preservation.
Index Terms:
@@
Human-in-the-Loop, Visual Navigation, Map Representation, Graph Representation, Gaze Behavior Generation, Implicit Behavior Cloning, Landmark, Dataset
I Introduction
@@
Robotic navigation remains a research problem of interest as advances in AI make intelligent mobile agents, such as delivery bots, more feasible. Many approaches have been proposed from SLAM [1, 2, 3], to graph-based methods [4, 5, 6, 7], to reinforcement learning (RL) [8] to enable robots to effectively move around a space in a goal-oriented manner. However, these methods suffer from high computational complexity, high memory usage, and low sample efficiency respectively. Visual navigation [9, 10, 11] methods circumvent several of these challenges by using neural networks to implicitly learn scene geometry and an optimal navigation policy without necessitating direct environment interaction.
@@
Figure 1: We propose a novel Landmark-Aware Visual Navigation (LAVN) dataset for map representation research. LAVN provides unique landmark information (in red) in specific scenes to aid model navigation and exploration performance and map representation learning. In this figure, a human expert in Room A (in blue) provides a video-based training trajectory (in green lines) for navigation given a goal image (green box). Rooms are depicted in different colors and represented with nodes. Our dataset facilitates topological map representation learning by injecting landmarks into a robot’s observation.
@@
Typically, visual navigation methods are evaluated with the image goal task, where an agent is given a goal image of some target and tasked with finding and reaching it in the environment. The success of state of the art (SOTA) methods at this task comes down to an agent’s ability to efficiently explore its environment. Several works focus on explicitly learning optimal exploration policies [12, 13, 14, 15]. However, they either abstract exploration out of the problem all together by assuming that the environment has been seen by the agent prior to test time [10], with the hope that their networks will implicitly learn an adequate exploration strategy from the data due to the specific task structure [16], or explicitly define the exploration strategy for the agent by hard coding it into the architecture [11].
Of those methods that learn to explore, several niches have arisen such as adding an exploration term to RL reward functions with [13] and without [17] behavioral cloning, building metric maps or graphs with [14] and without [18] semantics, and learning to propose exploration waypoints that provide helpful subgoals during navigation for local policies [15]. These methods suffer from similar downsides as those mentioned previously with the added potential for also requiring full ground truth semantic labels of an environment, which can be extremely costly to acquire. Instead, a much simpler approach would be to explicitly learn good exploration techniques in a human-centric manner. Learning human-centric exploration strategies in a supervised manner would allow for much simpler networks to effectively solve navigation problems, thus reducing the computational complexity, high memory usage, and low sample efficiency issues of SOTA. However, there is very limited data that would allow for this kind of learning.
In this work, we present a Landmark-Aware Visual Navigation (LAVN) dataset of human-guided environment exploration videos collected from the photorealistic Habitat virtual environment as well as several real world environments. We collect RGB image observations and human point click pairs as a human guides an agent in exploring each environment. These point clicks represent subgoals that guide the agent incrementally though the space and can be used for supervised learning of human-centric waypoints for environment exploration. Additionally, we expect that having landmarks when exploring has the potential to improve current SOTA performance on the map-building and localization tasks inherent in navigation. To this end, the human annotator also provides examples for helpful navigational landmarks, which will allow for supervised training of a landmark prediction network as well. In this way, it becomes possible to directly train a network to explore an environment in a human-centric manner under multiple SOTA paradigms as demonstrated by the recent work [19, 20] in the visual navigation task. In addition, learning from such waypoint and landmark annotations allows for aligning the cognitive mindset of human navigation with mobile robots, enabling intuitive interactions with people to convey mobile robots’ intent in navigation via gaze behavior generation [21, 22, 23, 24, 25].
This paper is organized as follows: we summarize our dataset’s key differences from existing works in Section II. Then we present the details of data acquisition both in the virtual and real-world environments in Section III and IV, respectively. We conclude our work in Section V.
II Related Work
II-A Map Representation for Navigation
The success of navigation relies on an effective learning and understanding of the environment. Two main approaches for this are map representation learning in geometry-based [1, 2, 3] and learning-based [12, 8, 10, 26, 27] domains. However, the former requires careful and robust hand-engineered feature design and requires intensive computation (e.g. keyframes in SLAM). The advancements of Deep RL in the second category are partially attributed to the emergence of simulators designed for photorealistic simulation, including Habitat AI [28, 29] with datasets Gibson [30] and Matterport [31]. While these simulators enable large scale model training, they are also expensive to build, can run slowly enough to mitigate their ability to produce large numbers of training samples, and introduce the sim2real problem when deploying models in the real world. To combat this, methods like NRNS [11] demonstrate the effectiveness of using expert videos to train robots without interacting with environments.
II-B Human-Robot Interaction Datasets
There is a large body of datasets or benchmarks for Human-Robot Interaction, ranging from industrial [32] to social [33, 34, 35, 36, 37, 38, 39, 40] environments. As one of the fundamental tasks for robots, researchers have collected datasets for learning visual navigation with socially compliant behaviors [33, 34, 35, 37, 38, 39, 40, 21, 23], including intuitively conveying mobile robot’s navigation intent [22, 24] to humans. Apart from learning from other robots collectively [41], robotic visual navigation can benefit from human feedback via visual cues [42], language [43], gesture [44, 25], multimodal instructions [38, 45], or demonstrations [46].
Our work extends the line of efficient map representation using topological graphs [7, 6]. Similar to how a human [47] can quickly learn to navigate in a new environment by implicitly treating landmarks as nodes in their mental graph [47], we hypothesize that a robot can acquire the ability to navigate in novel places efficiently by intelligently injecting landmarks into its own topological graph. However, little attention has been paid to such datasets, resulting in a lack of data that includes landmark information for this research in the literature. Our dataset hereby aims to fill this gap by providing a series of human-expert-annotated videos with landmarks both in virtual and real-world environments, including indoor and outdoor scenes. We summarize the key characteristics of recent datasets in Table II-B.
TABLE I: Summary of dataset characteristics for visual navigation. We provide a unique dataset with beneficial navigational landmark labels in human trajectory demonstrations in both real world and simulated environments. Add.: Additional; Mod: Modality; Sem: Semantics; LM: Landmark.
Dataset Human Traj. Demo Sim Real Small Scale No Add. Mod. No Sem. LM
Habitat-Web [46]
RUGD [42]
RealEstate20k [48]
EmbodiedQA [49]
SCAND [33]
@@ II-B Human-Robot Interaction Datasets
Figure 2: Samples of RGB images in our dataset captured from an ego-centric camera. Video frames start from left to right. Landmarks are visualized with red dots. Samples in the virtual dataset (1st and 2nd rows in blue) are visually realistic compared to the real-world observations (3rd and 4th rows in green), which makes landmark data acquisition scalable to serve real-world research purposes.
III Virtual Dataset
@@
The main objective of the data collection is to provide human navigation demonstrations for robots in a real-world environment. As video has been proven to be an effective way to teach robots to accomplish different tasks [11], we hereby collect trajectories in video format. Below, we detail the process of data collection for the virtual dataset.
III-A Environments
@@
We provide human-click annotations for a selection of the Matterport [31] and Gibson environments [30] in Habitat-Sim [28, 29] to facilitate virtual environment exploration. We provide a full list of annotated environments with the data. Habitat is already used to train many visual navigation methods, so having in-domain, landmark-aware data will allow for direct extensions of these methods for greater ease of model comparison. Additionally, it allows for greater data collection volume and lower data processing time.
We aim at randomly selecting rooms evenly distributed by room IDs to ensure diversity in room types and sizes, as detailed in the Appendix.
III-B Data Collection and Representation
@@
To start, the human navigator is placed at a random, navigable point in an environment, and tasked with visiting as many rooms as possible while identifying distinct landmarks in each room. The current first person observation of the environment is shown on a screen, and the human moves through the environment by left clicking on this image towards the direction they would like to move. There are three discrete moving action choices: moving forward 0.25m, turning left $15^{\circ}$, and turning right $15^{\circ}$. To specify a landmark, the navigator right clicks on the observation image over the object or region of space they choose. We provide a human point-click supervisory signal for each time step of the trajectory, but these signals could easily be masked to generate data to train temporally abstracted waypoint predictors.
@@
A graph is built as the navigator moves through the space, where nodes are added each time an action is taken. Each node contains one RGB observation, one depth image, the human navigational point click, and the ground truth odometry (including ground truth location and orientation). Edges are drawn between consecutive nodes only, however it is possible to perform loop closure on the graph using the odometry provided. Each graph also contains a sequential list of the actions taken to move through the environment. A landmark is defined as a unique and distinct object, distinguishable from other common objects in the scene. A list of pairs of landmark point-click image coordinates and their corresponding nodes is also provided. Each human-guided trajectory is terminated after 500 actions or when the annotator has navigated through the entirety of the space, whichever comes first.
@@
Figure 3: Samples of RGB images in the real world environments with landmarks annotated by red dots. Trajectories are recorded in comprehensive environments both indoors and outdoors.
IV Real World Dataset
IV-A Environments
@@
IV-B Data Collection and Representation
@@
In order to collect consistent data between the real-world data and simulation, the human annotator is asked to follow the same protocol as in the virtual environment data collection described in Section III-B. During navigation, the annotator holds a camera (iPhone13Mini [50]’s built-in camera) naturally around 1.5m high off the ground facing forwards. Unlike in simulation where agents can move with exact distances or turning angles, real-world environments require that each action taken by the human navigator is inherently inexact. Therefore, the navigator is instructed to follow the same discrete action choices by their own judgement. Specifically, the expert takes a picture (saved in JPEG) in every step (around 0.7m), or in every left or right turn (roughly $15^{\circ}$). This also presents the real-world data with complex noise or perturbations, which can improve downstream method robustness.
(a) Frames
(b) LM
Figure 4: Stacked histograms depicting ##\##Frames and ##\##LM (landmarks) per trajectory. (a) The majority of trajectories consists of the maximum ##\##Frames 500. (b) Most trajectories in the Gibson rooms consists of only a small number of annotated landmarks (e.g. $<25$), while it increases in Matterport due to its larger and more complicated environments.
TABLE II: Statistics of the LAVN dataset.
##\##Traj. Total Total ##\##Virt. ##\##Real World
##\##Frames ##\##Landmarks Envs Envs
310 103,998 14,281 300 10
@@
Resolution in the original images was 4032x3024. To promote consistency between the virtual and real-world data, the images are resized to be the same resolution of 640x480 as in the simulation and saved in JPEG format with quality $85\%$. Landmarks are annotated offline after the videos are saved. Samples from our dataset are visualized in Fig. 2. Both virtual and real-world images are photorealistic and can benefit downstream research in the real world. Also notice that trajectories are recorded in ten diverse scenes in both indoor and outdoor environments shown in Fig. 8. Our proposed dataset provides a unique benchmark for both learning and evaluation to serve the visual navigation community. Our dataset statistics are summarized in Table II and Fig. 4. Observe that most trajectories reach the maximum number of frames (500). On the other hand, a wide variety of trajectory horizon lengths (#$\#Frames/Traj.$) and number of landmarks (#$\#LM/Traj.$) is covered which benefit robot learning in both short [51] and long horizon [52, 53, 54] tasks.
(a) Action Stats
(b) Waypoint Density Map
Figure 5: (a) Stacked bar chart of actions in virtual and real-world environments. (b) Waypoints heatmap over the entire dataset.
For both datasets, the graph is saved as a json file. We detail the file organization and examples of dataset usage in the Appendix.
Real-world Relevance. The photorealistic images from virtual datasets (Gibson and Matterport), combined with our real dataset, support robot visual navigation learning tailored for real-world deployment.
V Conclusion